The first woman is 8 years younger than a man, and the man is 9 years older than the second woman. If the sum of their ages is 300, what is the age of the first woman?

Answer by stanbon(75887)

The first woman is 8 years younger than a man, and the man is 9 years older than the second woman. If the sum of their ages is 300, what is the age of the first woman?
------------------------
1st = m - 8
m = 2nd + 9
-----
m + 1st + 2nd = 300
----
m + m-8 + m-9 = 300
-------------------------------
3m -17 = 300
------
3m = 317
m = 15yrs 8mths
-----
Ans: 1st = m-8 = 7 yrs 8 mths
-------------------------
Cheers,
Stan H.
--------------------